# The evolution of inharmonicity and noisiness in contemporary popular music

### v1

@@ 4.5 HR-inharmonicity and peak prominence, weighted audio
This ends the core part of the article. The following sections elaborate on several elements from sections 3 and 4.
5 On the influence of technology on music
The example shown in section 2, Figure 1, features audio produced by a composite setup of digital instruments: superimposition of a Spectrasonics Omnisphere bass with a Eurorack setup involving FM synthesis, voltage-controlled amplifier, envelope, and ring modulator modules. It would have been difficult to produce using instruments based on the principle of the acoustic resonance. The present section elaborates on the relationship between technology and the resulting music. In section 5.1, we mention relations between the resources of the performance and the musical output. In section 5.2, we briefly focus on modern music production technology. In section 5.3, we explain the reasoning that brought us to study inharmonicity and noise in the first place. In particular, greater inharmonicity may be a case of CPM at least partially abandoning musical notes as a building block.
5.1 Resources of the performance and musical idioms
Music involves what Parry, (1911) calls the resources of the performance, such as the human voice and musical instruments. In Western classical music, a widespread class of such resources involves instruments based on acoustic resonators, of which examples are the vibrating string and the vibrating air column. Such acoustic resonators produce waveforms that feature a strongly periodic behaviour. The periodic part of such waveforms is usually modeled as the weighted sum of components based on a fundamental frequency and a series of overtones, whose frequencies are (close to) integer multiples of that fundamental (Young,, 1952).
According to Huron and Berec, (2009, p. 103), idiomaticism is, ‘of all the ways a given musical goal or effect may be achieved, the method employed by the composer/musician [that] is one of the least difficult’. Given an acoustic instrument that involves an acoustic resonator, it is easy to produce a harmonic or quasi-harmonic sound. Harmonic or quasi-harmonic sounds may therefore be described as idiomatic to instruments involving acoustic resonators.
The resulting sounds are used as elements of the musical discourse. In the case of acoustic resonators, the pitch impression derived from the harmonic or quasi-harmonic sound is denoted as a musical note, which is then used as a building block for writing musical phrases and chords. In Pascall,’s (2001) terminology, the idiom resides in the possibilities (for instance, the musical phrases and the chords) offered by the characteristic elements stemming from the use of the instrument. The properties of the sounds produced by acoustic resonators may also condition specific aspects of the musical discourse. For instance, the position of harmonics is an important factor in the determination of the musical temperament (Barbour,, 2004), and rules of Western counterpoint, such as the prohibition of parallel unisons, octaves, and fifths, can be linked to the properties of the waveform resulting from acoustic resonators (Huron,, 2001).
By providing a larger variety of resources, technological progress may influence musical idioms. In the 20th century, one outcome of technological progress was signal analysis. Musicians such as Gérard Grisey took advantage of technology to produce spectral music, in which the score is derived from the positions of the harmonics of an acoustic instrument. One musical parameter becomes the deviation from the positions of the harmonics, resulting in a gradation from a harmonic texture to an inharmonic texture (Rose,, 1996). In this case, the idiom according to Huron and Berec, (2009) is the orchestral texture, and the idiom according to Pascall, (2001) is the gradation.
More generally, the 20th and 21st centuries brought considerable technical innovation in music technology (Webster,, 2002). Examples include non-acoustic musical instruments such as the Ondes Martenot (1928) (Orton,, 2001), the Moog analog synthesiser (1964) (Porcaro,, 2001) and digital synthesisers such as the FM-based Yamaha DX7 (1983) (Mattis,, 2001), which played a key role in popular music from the ’80s (Lavengood,, 2017, p. 36). Another key innovation resides in the practice of sound recording, which has been in a constant state of development since 1877 (Mumma et al.,, 2001). At least some technologies from the 20th and 21st centuries can be associated with musical idioms. For instance, continuously evolving frequencies can be identified as idiomatic (in the sense of Huron and Berec,, 2009) to the Ondes Martenot. Distortion can be identified as idiomatic to the use of analog circuitry, resulting in an idiom (still in the sense of Huron and Berec,) that is a definitive feature of musical genres such as heavy metal.
5.2 The ‘utopian vision of a boundless space for musical exploration’
The advent of specific idioms is not the only consequence of technological innovation. Music-related technology can assume various forms and is involved in various stages of the processes of producing music and engaging with it. One of these aspects is the compositional process itself. For instance, before the 16th century, paper was expensive and erasable pencils did not exist (Charlton and Whitney,, 2001). As a result, composers did not have the opportunity to sketch and draft musical ideas (Bent,, 2001). The introduction of cheaper paper and erasable pencils modified the composition process. More recently, one consequence of technological progress in the composition process is the use of the recording studio for its own creative potential (Moylan,, 2014). Spicer, (2004) directly links ‘accumulative’ and ‘cumulative’ forms in pop-rock music to the rapid advances in recording technology.
During the 1990s, with the advent of the home studio, the practices that were previously exclusive to the professional studio became universally available. Almost anybody with a computer could now assume a musician–engineer hybrid role (Pras et al.,, 2013; Bell,, 2014) in a music production process to which ‘[t]he micro-manipulation of digital audio has become more and more the primary focus’ (Théberge,, 2001).
We have argued above that the recourse to musical notes can be considered as idiomatic to the use of acoustic resonators. In other words, instruments involving acoustic resonators favor the constraint (McPherson and Tahıroğlu,, 2020) according to which music should involve symbols referred to as ‘notes’, each of which denotes a harmonic or quasi-harmonic complex tone based on a particular fundamental frequency that remains approximately constant throughout the note’s duration. In the studio environment, this particular constraint disappears. One important issue becomes the determination of the idiom used by musicians in the absence of such a constraint. The issue relates to McPherson and Tahıroğlu,’s (2020) concern, according to which ‘even if we were to achieve the utopian vision of a boundless space for musical exploration, we would still be left with the question of what possibilities musicians would choose to explore within it’.
5.3 To what extent do idioms of popular music involve musical notes?
Musique concrète is an instance of music in which musicians have abandoned the musical note as a building block while working in the recording studio (Schaeffer,, 2020). This music genre, derived from Schaeffer,’s (1966) experiments, often does not involve elements that prompt a clear sensation of pitch. Still, musique concrète is not entirely devoid of pitch. According to Yost, (2009), ‘[m]usic without pitch would be drumbeats, speech without pitch processing would be whispers, and identifying sound sources without using pitch would be severely limited’. Such a description does not apply to musique concrète, as evidenced by pieces such as Temps de Pointe in Pierre Henry’s Mouvement-Rythme-Etude or Bernard Parmeggiani’s Ondes. Nevertheless, transcribing musique concrète to standard staff notation is generally difficult at best. Instead, transcriptions of musique concrète have been seen to involve freely drawn shapes that may reflect dynamics or timbre perception (Favreau et al.,, 2010).
From a commercial or large-scale social perspective, musique concrète might be considered a niche. The music charts indicate that the music that is most listened to is recent popular music. Popular music is a broad category that is distinguished from folk music and art music by its use of recorded sound as the main mode of transmission (Tagg,, 1982; Middleton,, 1990; Mazzanti,, 2019). It is characterized by commercial/industrial interests, entertainment, and strong ties to mass media (Frith,, 2004).
Deruty et al., (2022) use the term Contemporary Popular Music (CPM) to refer to popular music that is ‘contemporary’ in the sense of ‘current’ or ‘present-time’. CPM is known for its focus on technological innovation and cross-cultural and cross-genre influences (Mazzanti,, 2019; Frith,, 2004). Examples of CPM genres include post-rock, rap/hip-hop, electronica, and non-Western genres such as K-pop, J-pop, Bollywood, and reggaeton. Bertin-Mahieux et al., (2011) use the term contemporary popular music in relation to tracks released from 1922 onwards. Here, we use the term to refer to popular music released since the beginning of the 1970s. A milestone for the start of this period could be Sgt. Pepper’s Lonely Hearts Club Band, as the first release of The Beatles’ studio era (McCartney and Miles,, 1997). The dataset that we explore includes popular music released since 1961. Like musique concrète, recent popular music is produced in the recording studio. In Parry,’s (1911) terms, these two apparently very different categories there share the same ‘resources of the performance’ (the studio), and therefore have similar constraints and affordances.
Has popular music abandoned musical notes as its building blocks? Huang et al., (2020) distinguish several building blocks in the making of popular music, amongst which are melody, harmony, and percussion. The ‘melody’ and ‘harmony’ building blocks suggest that popular music can be at least partially transcribed to musical notes, either automatically (e.g., Bertin-Mahieux et al.,, 2011) or manually (e.g., Tagg,, 1982). The ‘percussion’ building block has been considered important enough to form the subject of research in both automatic transcription (Wu et al.,, 2018) and musicology (e.g., Mowitt,, 2002). From the perspective of physics, percussion instruments such as snare drums have been identified as featuring clear stochastic noise components (Fletcher and Rossing,, 2012, pp. 602-606). From the perspective of signal processing, percussive elements have been considered noisy impulses, forming ‘vertical ridges with a broadband frequency response’ (Fitzgerald,, 2010). Both perspectives suggest that the transcription of percussion to musical notes is, at best, difficult.
Percussion is not the only aspect of CPM that creates difficulties for meaningful transcription to musical notes. For instance, the practice of sampling has been documented to bring real-life noisy sounds into the music (Forman and Neal,, 2004, p. 408). The noise surrounding each partial in guitar distortion (Berger and Fales,, 2005, p. 184) may blur frequency values and therefore make pitch identification difficult. The fact that not all popular music can be meaningfully transcribed to notes (i.e., staff notation) raises the following question: is it possible to quantify the extent to which popular music deviates from being the result of a combination of ‘notes’? In other words, to what extent does it deviate from being produced (or producible) by some combination of acoustic resonators or technological devices that emulate acoustic resonators? We hope this article provides the beginning of an answer to this question.
6 A study of HarmonicRatio
In this section, we shall look more closely into the HarmonicRatio feature introduced in section 3.2.1. While it is not the focus of this paper to provide a complete understanding of the relations between the notions of ‘harmonicity’, ‘inharmonicity’, and the HarmonicRatio feature, we need an intuitive understanding of the feature’s behaviour, so that we can interpret the results of the subsequent measures involving HarmonicRatio.
6.1 Conventions

### v2

@@ 3.3.5 HR-inharmonicity and peak prominence, weighted audio
3.4 Relations with some other diachronic studies
Mauch et al., (2015) investigate the US Billboard Hot 100 between 1960 and 2010. Using music information retrieval and text-mining tools, they demonstrate quantitative trends in their harmonic and timbral properties. Using NNLS Chromas (Mauch and Dixon,, 2010) and MFCCs, they find that ‘1964, 1983 and 1991 are periods of particularly rapid musical change’. They remark that ‘other measures may give different results’, which is indeed the case in this study, where we observe the fastest changes occurring throughout the 1960s (see Figures 5 and 7), as well as turning points in 1972, 1986 and 2000. The features used in Mauch et al., (2015) do not seem well-correlated to the features we use, suggesting that chromas and MFCC may fail to recognise certain properties of the music. As a general rule, it may be that the interpretation of feature-based studies should be limited to the interpretation of the features themselves rather than to ‘the music’ as a whole.
In Figure 5, PC1 and PC2 are orthogonal. Therefore, the variance for the 2D data can be estimated as the sum of the variances along each axis. Figure 9 represents the sum of the variances along PC1 and PC2 for each dataset, as well as for the year-by-year data of the BEA dataset. The results are identical in terms of comparison for the original and PC representations. The following points are worth noting.
  1. 1.
    The variance of popular music in terms of the features we use is higher than that of orchestral music after 1974, and higher overall. It suggests that in terms of noise and HR-inharmonicity, popular music uses a wider ‘space for musical exploration’ than orchestral music.
  2. 2.
    The variance of musique concrète in terms of the features we use is greater than that of popular music. In terms of noise and HR-inharmonicity, Musique concrète takes greater advantage of the lack of constraints in the production process.
  3. 3.
3.5 Influence of the medium and remastering processes
The medium on which the music is recorded may contribute to noise in the recording. According to Brandt et al., (2019), early recording media, such as wax cylinders, had signal-to-noise ratios (SNRs) of below 40 dB. Vinyl discs have SNRs of 55 to 60 dB, and magnetic tape storage SNRs of 60 to 70dB. To evaluate the extent to which the medium contributes to noisiness, we compute the difference between the noisiness in the original BEA files and the same files on which we add white noise corresponding to an SNR of 40dB. The median of the absolute value of the difference is 0.000007. In Figure 4, typical noisiness values for the BEA dataset are shown to be approximately $0.3$. The comparison suggests that the contribution of the medium to the total amount of noise is negligible.
Music albums may be remastered. Deruty, (2011) suggests that remastering significantly affects the album’s loudness. We compared HR-inharmonicity values for the songs from the original and remastered version of the British band, The Cure, as Deruty, (2011) did for loudness values. The median of the absolute value of the difference is 0.015 for non-weighted audio and 0.008 for weighted audio. In Figure 3, typical HR-inharmonicity values for the BEA dataset range between 0.3 and 0.63. Along with the influence of the background noise, the comparison suggests that the study’s results are robust to remastering processes.
4 Analysis of specific tracks and artists
In this section, we focus on the music of particular artists using the analysis of weighted audio proposed in Section 3.3.5. We suggest links between the analysis results and production methods.
4.1 Selected artists
Figure 10 illustrates noisiness and HR-inharmonicity values for several artists and speech tracks (weighted audio). Artists were chosen that lie on the edges of the distribution so that it is easier to understand the perceptual meaning of the two dimensions. Figure 11 shows the same data after PCA. We go through each element in Figures 10 and 11 so as to identify links between the feature values and aspects of the corresponding audio content.
  1. 1.
    Initial reference: speech tracks. The speech tracks in Figures 10 and 11 (white dots) are ‘interludes’ or ‘skits’ as found in hip-hop music. They generally have a low PC2 value. Low PC2 values for speech tracks illustrate how a sound can be non-inharmonic while not featuring stable pitch values. Higher PC2 values are observed in some of the examples of distorted voices. Higher PC1 values are observed in the examples that feature a high background noise.
  2. 2.
    Figure 10: BEA dataset, normalised noisiness and HR-inharmonicity corresponding to particular artists as well as to monodic speech tracks. For each artist, the contours represent the isolines at half-height. Contours are evaluated on a mean filter of the distribution for better readability. The artist’s name is centered on the median. The white dots correspond to individual monodic and possibly processed speech tracks (the ‘skits’). The red and blue stars denote the feature distributions for the orchestral and piano music datasets, as in Figure 6.
  3. 3.
    Exemplification of the 1961–1986 evolution. The combined evolution of PC1 and PC2 from 1961 to 1986 is exemplified by the productions of Barbara and My Bloody Valentine. My Bloody Valentine’s music is characterised by complex guitar textures with open tuning (Leonard,, 2021), pitch bending, tremolo (Di Perna,, 1992), and an extensive effects rig (Double,, 1992). Although chords can be identified, individual note perception is difficult, consistent with (a) the high number of inharmonic relations in their tracks (top right of the PC1–PC2 representation) and (b) the idea that pitch perception in inharmonic sounds is more ambiguous than in harmonic ones (Schneider,, 2000; Schneider and Frieler,, 2009). According to music critic Anthony Fantano, the guitar in their album ‘Loveless’ is ‘slathered with waves of pink smog’, obscuring and transforming the perception of the guitar (Fantano,, 2013). The examples of Barbara and My Bloody Valentine suggest that, between 1961 and 1986, recorded music shifted from acoustic, tonal instruments playing content that could be transcribed to a score, to heavily processed, noisy, and inharmonic studio-made content, including drums.
    Figure 11: Same data as in Figure 10, PCA.
  4. 4.
    Noise from distortion. The position of Dead Kennedys in Figures 10 and 11 shows that they exhibit very high total noise and inharmonicity, with the proportion of inharmonicity relative to noise being around the median for that level of noise. As a punk rock band from the 1980s, Dead Kennedys’ music includes noisy vocals, drums, bass, and distorted guitar. The noise likely stems from the vocals, drums, and guitar. Drums cover a wide frequency range, and in heavy-metal-style distortion, noise surrounds each harmonic rather than layering with them (Berger and Fales,, 2005, p. 184). Figure 12 illustrates that the noise in their music is spread across different frequencies. Their PC2 values are lower than My Bloody Valentine’s, likely due to the lower volume of guitar parts in the mix and fewer apparent layers.
  5. 5.
    Exemplification of the 1986–2020 evolution. The combined evolution of PC1 and PC2 from 1986 to 2020 is illustrated by the productions of Dead Kennedys and London Grammar. Both bands have an inharmonicity-to-noise ratio near the BEA dataset median, but the total amount of noise and inharmonicity is very low for London Grammar and very high for Dead Kennedys. London Grammar, an indie pop band from the 2010s, creates studio- and synth-oriented music (MusicTech,, 2023), with carefully produced, atmospheric tracks (Senior,, 2014; 7Digital United States,, 2013). Their music features lead vocals with smoothed plosives, noiseless yet rich multi-layered arrangements, and heavy reverb. These layers and reverb likely explain the higher PC2 values compared to Joanna Newsom’s music. The examples of Dead Kennedys and London Grammar suggest that, between 1986 and 2020, recorded music shifted from simpler, noisier tones to more polished, multi-layered studio productions. However, this evolution is less pronounced than the shift observed between 1961 and 1986.
  6. 6.
    Slightly inharmonic music based on acoustic instruments. Joanna Newsom’s music has a higher PC2 value than Barbara’s, but lower total noise and inharmonicity, indicating a higher inharmonicity-to-noise ratio in Newsom’s music. This ratio stems from inharmonic tones. Newsom’s music, featuring fewer layers than London Grammar’s, suggests that fewer elements may result in less inharmonicity (see Appendix A.1.3 for further discussion). Her music consists mainly of vocals accompanied by harp, with occasional piano and orchestra. The harp’s difficulty in tuning (Cathcart,, 2018) and its initial ‘twang’ with higher frequencies converging to the pitch (Fletcher,, 2000) may explain the higher PC2 values. The inharmonicity of piano strings (Rasch and Heetvelt,, 1985) also aligns Newsom’s music with piano dataset values. However, the PC2 value remains below the median, meaning the inharmonicity relative to noise is lower than average. Low noisiness likely results from clean instrumentation, minimal use of drums, and the absence of prominent plosives in the vocals.
  7. 7.
    Noise from sampling. Public Enemy’s music is as noisy as Dead Kennedys’ but less inharmonic. The two Public Enemy albums in the dataset, released in 1988 and 1990, are characterised by the recombination of numerous samples across multiple layers (McLeod and DiCola,, 2011, pp. 22-26). The sound has been described as ‘part musique concrète, […] a noisy collage of sputtering Uzis, wailing sirens, fragments of radio and TV commentary […], all riding on rhythms articulated by constantly changing drum voices […], off-kilter loops, aliased or scratchy samples, and high-pitched spiraling sounds’ (Forman and Neal,, 2004, p. 408), explaining the high noisiness. The lower inharmonicity-to-noise ratio compared to Dead Kennedys may stem from Public Enemy using fewer pitched elements. Like Joanna Newsom’s music, Public Enemy’s music has a lower-than-average inharmonicity-to-noise ratio (PC2). However, unlike Newsom’s music, the total amount of inharmonicity and noise in Public Enemy’s music is very high (PC1).
  8. 8.
    Does studio work lead to more inharmonic partials? The National’s music is less noisy than My Bloody Valentine’s, with a similar proportion of inharmonicity deriving from partials. Like My Bloody Valentine, The National’s production involves extensive studio experimentation (Doyle,, 2017). The band has two guitarists who use large pedal-boards, focusing on ‘textural soundscaping’ rather than virtuosity. One of their goals is to create a ‘lattice work of notes’ (Guitar.com,, 2017). An example of this can be found in the album ‘Sleep Well Beast’, where ‘the pianos are playing off each other by an eighth note’ (Guitar.com,, 2017), increasing the number of simultaneous pitch values. Since more distinct tones result in greater inharmonicity, such techniques likely contribute to higher inharmonicity values. Additionally, much sustain and reverb will also lead to increased inharmonicity because of tones that are close together in pitch in the same part or voice overlapping in time.
4.2 Possible causes for high PC1 and PC2 values
PC1 values are generally higher in the case of popular music than in the case of orchestral music (see Figures 6 and 8, as well as Figure 13 below). Possible factors for higher PC1 values may involve loud drums (Dead Kennedys, Public Enemy), distortion (Dead Kennedys, My Bloody Valentine), noisy samples (Public Enemy), and vocals with loud plosives (Dead Kennedys, Public Enemy).
4.3 High inharmonicity—to what extent?
Given that we perform the analysis on final stereo tracks, it is difficult to verify this conclusion in the general case. However, it is possible to analyse a particular example to understand one way to reach high PC2 values. The song ‘Sometimes’ by My Bloody Valentine features particularly high PC2 values. At the end of the song, a solo guitar chord corresponds to even higher PC2 values (My Bloody Valentine,, 1991). We isolate this part and evaluate its power spectrum. Figure 14, top, shows that the part is a quasi-harmonic 37.5Hz complex tone with its fundamental missing. Figure 14, middle, shows that the relative loudness of the overtones is high, which explains why this single complex tone was initially perceived as a chord: we hear some of the overtones as independent notes. Figure 14, bottom, shows the frequency difference between consecutive partials. The difference is not constant, which makes the complex tone strongly inharmonic.
Figure 14: Excerpt from My Bloody Valentine, ‘Sometimes’, from the album ‘Loveless’. The excerpt’s timing is 5’06 to 5’19. (a) Power spectrum. (b) Power spectrum weighted with ISO226:2003, 50 phon. (c) Frequency difference between consecutive partials. The mean frequency difference (37.5Hz) is drawn in the top and middle diagram as the blue dotted line.
Such properties of the signal are not specific to this particular song. For instance, we can witness a similar tone architecture in the keyboard part at the end of Alt-J’s ‘Hunger Of The Pine’ (Alt-J,, 2014)—albeit with fewer inharmonic overtones. It is also worth noting that ‘Loveless’ has been highly influential, being an inspiration to artists such as The Verve, Oasis, Deerhunter, M83, DIIV, Deafheaven and Coldplay (Hudson,, 2021). It is rated by BestEverAlbums.com as the second best album of 1991, second only to the extremely successful ‘Nevermind’ by Nirvana.
Schneider and Frieler, (2009) studied inharmonicity in bells from the belfry of Brugge, noting that the SPINET pitch detection algorithm (Cohen et al.,, 1995) fails to yield interpretable results with single bell samples. They also observed that the use of bells ‘hampers [human] detection of the pitches’, complicating voice following and understanding of tonal functions. This raises the question of whether inharmonicity in CPM similarly obstructs algorithmic pitch detection, voice tracking, and tonal function comprehension.
We compared PC1 and PC2 values for bells from Dutch belfries [^3] https://revivethis.org/sample-pack-bells/ with those for the BEA dataset. Figure 15 shows the results, indicating that bells generally exhibit higher PC1 and PC2 values than popular music. However, the values for the complex tone in Figure 14 (My Bloody Valentine, ‘Sometimes’) exceed the median bell values in both PC1 and PC2. This complex tone reflects a higher total noise and inharmonicity (PC1) and a higher proportion of inharmonicity from intervals between partials (PC2) than most bell sounds. It suggests that, in CPM, some instruments (in this case, the electric guitar, which is not a marginal instrument) may introduce enough inharmonic relations between partials to hinder human and algorithmic pitch recognition.
5 Conclusions
The initial question, as introduced in Section 1 and elaborated upon in Section 2.2.3, can be phrased as follows: given that recent popular music, unlike Western classical music, is not subject to the constraint of (almost exclusively) using acoustic resonators (as employed in acoustic instruments), to what extent do the two music categories differ in terms of properties that can be related to the use of acoustic resonators? The short answer to this question is that recent popular music is both noisier and more inharmonic than Western classical music (Section 3.3, Figures 6 and 8). It is also more inharmonic than musique concrète and slightly less noisy.
We now summarise the methods on which the study is based.
  1. 1.
    We directly derived the HR-inharmonicity feature (Section 3.2.1, detailed in Appendix A.1) from the MPEG-7’s HarmonicRatio feature, which in turn has been used to measure pitch strength.
  2. 2.
    We used the noisiness feature (Section 3.2.2, detailed in Appendix A.2), derived from the MPEG-7’s AudioFlatnessType feature, primarily characterised as being the opposite of peak prominence. Noisiness denotes a locally ‘flat’ spectrum.
  3. 3.
    We argued that, as the human ear is not equally sensitive to all frequencies, weighting the audio using equal loudness contours prior to feature evaluation can provide results that better reflect what a human listener actually hears (Section 3.2.4, detailed in Appendix B). Measures showed that the weighting increases the proximity of perceptually similar tracks and reduces the number of outliers.
  4. 4.
    As HR-inharmonicity and noisiness are not independent of each other (Section 3.2.3, detailed in Appendix A.1.6), we used PCA to define two new features, PC1 and PC2. PC2 is particularly significant, as it measures the proportion of the total inharmonicity that is due to inharmonic relations between partials (as opposed to noise). It follows that it also measures the relative effect on pitch strength of inharmonic relations between partials and noise.
We now summarise the main findings of the study.
  1. 1.
    We showed that HR-inharmonicity of popular music is comparable to that of musique concrète (Section 3.3.1). At the beginning of the 1960s, it was similar to that of Western orchestral music, and it then increased to reach a peak in the mid-1980s. In Appendix A.1, we show that high HR-inharmonicity values, and, therefore, low pitch strength, may derive from noise and/or interaction between inharmonically related partials within and between complex tones. In the latter case, the degree of inharmonicity depends on both the proximity and energy of interacting partials. We also showed that higher HR-inharmonicity values prompt faster acoustic beating.
  2. 2.
    We showed that noisiness of popular music is comparable to that of Western orchestral music and lower than that of musique concrète (Section 3.3.2). It reached a maximum during the mid-1980s.
  3. 3.
    In regards to the frequencies to which humans are the most sensitive, popular music is generally slightly noisier and considerably more HR-inharmonic (lesser pitch strength) than orchestral music (Section 3.3.4). The total amount of HR-inharmonicity and noisiness combined (PC1) is greater on average in popular music than in Western classical music.
  4. 4.
    The proportion of HR-inharmonicity that is due to inharmonic relations between partials for popular music at the beginning of the 1960s was comparable to that of orchestral music (Section 3.3.4). It then increased rapidly until ca. 1972. Later evolutions are slower. It was therefore suggested that popular music released after the evolution forms a category of its own, which we refer to as ‘Contemporary Popular Music’ (CPM). The density of inharmonic relations between partials within CPM is, on average, significantly higher than that of music from other categories.
We linked feature values with examples drawn from the popular music dataset (Sections 4.1 and 4.2) and observed the following:
  1. 1.
  2. 2.
    HR-inharmonicity and PC2 are more difficult to pinpoint. We hypothesise that high HR-inharmonicity and PC2 values may derive from the use of ‘tonal’ sounds coupled with intensive use of the studio. Such use of the studio may be specific to CPM, as musique concrète, another category of studio-based music, is both noisier and less HR-inharmonic.
  1. 1.
    While high inharmonicity values may originate from both the intervals between complex tones and the inharmonicity of individual complex tones, increasing the inharmonicity of individual complex tones reduces the relative influence on inharmonicity of the intervals between complex tones (Appendix A.1.4).
  2. 2.
    According to Peeters, (2004, p. 20), ‘Spectral Flatness is a measure of the noisiness (flat, decorrelation)/sinusoidality of a spectrum’. According to (ISO,, 2001, p. 26), ‘[t]his descriptor expresses the deviation of the signal’s power spectrum over frequency from a flat shape (corresponding to a noise-like or an impulse-like signal). A high deviation from a flat shape may indicate the presence of tonal components’. According to the Essentia documentation, [^4] https://essentia.upf.edu/reference/streaming/_Flatness.html spectral flatness is ‘a measure of how noise-like a sound is, as opposed to being tone-like’. We demonstrated that Spectral Flatness fails to give high noisiness values for pink noise, undermining its purpose. We suggested filtering out the overall spectral envelope before evaluating spectral flatness (Appendix A.2).
  3. 3.
    In terms of noisiness and inharmonicity, we found no evidence of the ‘homogenization of the timbral palette’ reported by Serrà et al., (2012). On the contrary, the evolution of popular music towards CPM is accompanied by a significant increase in the diversity of noisiness and inharmonicity values, which evolves to be greater than that of orchestral music (Section 3.4).
  4. 4.
    The output of feature-based studies may be made more reliable by attenuating the influence of audio frequencies human listeners are less sensitive to (see Appendix B). In particular, the weighting function gives relatively much less weight to the lower frequencies. This suggests that it may be useful to weigh the initial signal using models such as ISO226:2003 (ISO,, 2003).
Finally, in Section 4.3, we compared the HR-inharmonicity and PC2 values in popular music with those for church bells. Although church bells generally involve more inharmonic relations between partials than popular music, we showed that there are instances in CPM of sounds in which there are more inharmonic relations between partials than most church bells. It has been suggested in the literature that inharmonicity in bells hampers both automatic pitch detection and human perception of pitch. We therefore expect similar difficulties to arise when detecting or perceiving pitch in passages from CPM songs where the values of HR-inharmonicity and PC2 are high. In such cases, transcribing the music into notes in staff notation may make little sense. Using the terms introduced in Sections 2.2.1 and 2.2.3, idioms of popular music may not always involve musical notes, even when the elements on which the idioms are based involve waveforms that feature a strongly periodic behaviour and convey an impression of pitch. This is a topic we intend to explore further in future studies.
Appendix A HR-inharmonicity and Noisiness
A.1 A study of HarmonicRatio
In this section, we examine the HarmonicRatio feature mentioned in Section 3.2.1. Though a full explanation of ‘harmonicity’, ‘inharmonicity’, and the HarmonicRatio feature is beyond this paper’s scope, an intuitive grasp of the feature’s behaviour is needed to interpret the subsequent HarmonicRatio measures.
A.1.1 Conventions